OpenStudy (anonymous):

the equation in standard form which passes through (3, –6) and (–2, –1) ?

OpenStudy (anonymous):

\[y=-x-3\]

OpenStudy (anonymous):

you need the method?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Not quite standard form, standard form would be x + y = 3.

OpenStudy (anonymous):

first find the slope via \[\frac{y_2-y_1}{x_2-x_1}=\frac{-6+1}{3+2}=\frac{-5}{5}=1\]

OpenStudy (anonymous):

typo \[\frac{-5}{5}=-1\]

OpenStudy (anonymous):

ehhh ?

OpenStudy (anonymous):

then use \[y-y_1=m(x-x_1)\] with \[y_1=-6, x_1=3,m=-1\]

OpenStudy (anonymous):

okay i got it (: